Please, am in dire need of the answer to this question. it is urgent.
A rectangular yard is to be constructed along the side of a house using a fence on three sides. The house wall will serve as the fourth side. Find the dimensions that produce the maximum area if 40 meters of fence is available.
Thanks.


Answer by checkley77(12844)

THE MAXIMUM AREA FOR A 3 SIDED FENCE HAS A RATIO OF 1:2:1.
THE SUM OF THESE RATIOS=4.
THUS WE DIVIDE THE LENGTH OF THE FENCING 40/4=10.
THUS THE SIDES ARE 10:20:10.
PROOF:
10*20=200 YD^2 ANSWER.
9*22=198 YD^2
11*18=198 YD^2.